OpenStudy (anonymous):

Find the x-intercept and y-intercept of the line. x-3y = -6 y -intercept: x -intercept:

OpenStudy (utterly_confuzzled):

To find the x-int plug 0 in for y To find the y-int plug 0 in for x

OpenStudy (calculusxy):

i recommend you to put this into slope-intercept form: y = mx + b -3y = -6 - x y = 2 + 1/3x y = 1/3x + 2

OpenStudy (calculusxy):

And then you can plug in 0's on the different variables depending on what you want to find. But use this info: \(\color{#0cbb34}{\text{Originally Posted by}}\) @Utterly_Confuzzled To find the x-int plug 0 in for y To find the y-int plug 0 in for x \(\color{#0cbb34}{\text{End of Quote}}\)